Tuesday, January 5, 2010

Letter from Crowley to USCG, ADEC and Alyeska/SERVS regarding Pathfinder grounding spill volume determination

Tug Pathfinder release of diesel fuel determined (a Crowley Press Release)

Crowley Situation Report #3 January 4, 2010 as of 1600

Busby Island Bay Response
Project Plan

SITREP #3- January 4, 2010 @ 1400 hrs

Pathfinder Status

The PATHFINDER is still moored at the Valdez Container Terminal (VCT). The tug is boomed and there has been no visible sheen since it was docked on December 29. There is 24 hr. security and spill response equipment has been provided by our oil spill removal organization Chadux.

Emerald Environmental Services has removed all of the remaining fuel oil and the tanks were vacuumed dry. In addition, the oil slops and dirty waste oil have been removed. The vessel engine room bilges have also been pumped dry.

On December 30 Ian Hogben, an independent Marine Surveyor, conducted a survey of the internal fuel and water tanks as well as the hull above the water line. The surveyor recommendations for towage of the tug will be provided in his written survey report to be provided this week. Based on the surveyor’s verbal recommendations the following work on the tug has been initiated:

• Global diving was contracted and they will arrive p.m. on January 4th. They will do temporary patching of the main water tank and tank 1 port as well as secure the openings to the keel coolers. Global also has the capability for high resolution underwater video if it is needed.
• Preparation for arresting the movement of the rudders.
• Activation of the reduction gear clutches for locking of the propeller shafts
• All fuel oil tanks have been secured and vents are being sealed.
• The after generator was converted to radiator cooling to eliminate the thru-hull water cooling line.
SERVS response gear and other operational equipment and supplies were transferred to the replacement tug Guardian. The tug Guardian was response ready at 2400 hrs on January 1, 2010. Additional electronic communication equipment will be transferred to the Guardian on January 5th.
The spill volume from the PATHFINDER grounding was finalized by the IMT environmental unit and notification provided to the Coast Guard and ADEC on January 4th.

Pathfinder Towage
Crowley is still in the process of preparing a towing plan for repositioning the PATHFINDER to a repair facility, location not yet determined. The tow plan will be approved by the United States Coast Guard prior to further movement of the vessel.

Monday, January 4, 2010

Updated ADEC Situation Report

DEC has updated its incident notification page for the Pathfinder grounding. Use this link to access updated situation reports throughout the response. A situation report for December 30, 2009 at 3:00pm is now available. Please note, the next situation report will be released from ADEC as the situation warrants one necessary.